# Deciding Local Unitary Equivalence of Graph States in Quasi-Polynomial Time

### v1

@@ Appendix F Computer-assisted study of 2-local complementation
  • $|V^{\prime}\setminus S^{\prime}|\leqslant|V\setminus\textup{{supp}}(S)|$
  • there exists no set $A\subseteq S^{\prime}$ such that $G\star^{2}S=G\star^{1}A$.
@@
The proof is constructive, in the sense that we construct $G^{\prime}$ and $S^{\prime}$ from $G$ and $S$. According to Proposition 2.3, there exists set $S_{1},S_{2}\subseteq\textup{{supp}}(S)$ such that $S_{2}$ is 2-incident and $G\star^{2}S=G\star^{2}S_{2}\star^{1}S_{1}$. Also, the existence of a set $A\subseteq S_{2}$ such that $G^{\prime}\star^{2}S_{2}=G^{\prime}\star^{1}A$ implies $G\star^{2}S=G\star^{1}(A\Delta S_{1})$. Let $G^{\prime}=G$ and $S^{\prime}=S_{2}$, then apply the following operations:
@@
    Remove the vertices of $V^{\prime}\setminus S^{\prime}$ of degree 0;
@@
It is easy to generate each graph of $\mathcal{G}_{k}$, although the number of elements in $\mathcal{G}_{k}$ grows double exponentially fast with $k$. $S$ can be defined as a list of words in $\{0,1\}^{k}$ of weight at least 2. More precisely each vertex of $S$ is uniquely associated with a set of $V\setminus S$ of size at least 2, its neighbourdhood. Furthermore, the 2-incidence of $S$ implies that $S$ is uniquely determined by the set of its vertices of degree at least 4. Indeed, starting from a set containing only vertices of degree at least 4, the conditions of the form "$S\bullet\Lambda_{G}^{K}=0\bmod 2^{r-k-\delta(k)}$" translate into a procedure to find which vertices of degree 3 then 2 need to be added to the set so that $S$ is 2-incident. This proves that there is a bijection between $\mathcal{G}_{k}$ and lists of words in $\{0,1\}^{k}$ of weight at least 4. Thus, the size of $\mathcal{G}_{k}$ is exactly given by the formula
@@
For $k=1,2,3,4,5$ and $6$, the size of $\mathcal{G}_{k}$ is respectively $641,1,1,2,2^{6}=$ 64, and $2^{22}\sim 4\times 10^{6}$ which is suitable for computation. But, even for $k$ as low as seven, the size of $\mathcal{G}_{7}$ is $2^{64}\sim 2\times 10^{19}$.
@@
According to Lemma F.1, this is enough to prove Lemma 4.1:
@@
Given a set $S$, we define the strict neighbourhood of $S$ in the graph $G$ as the set of vertices not in $S$ that are connected to at least one vertex in $S$: $\delta_{G}(S)=\{v\in V\setminus S\leavevmode\nobreak\ |\leavevmode\nobreak\ N_{G}(v)\cap S\neq\emptyset\}$. Let us prove, by induction on $|\delta_{G}(S)|$, that for any 2-incident independent set $S$ such that $S$ does not contain any twins and $|S|\leqslant 12$, $G\star^{2}S=G$.
First, by brute force, we check that the property is true whenever $|\delta(S)|\leqslant 6$ (see details in Appendix F, code available at [9]).
@@
Let $w\in\delta(S)\setminus\{u,v\}$, and let $G^{\prime}=G[V\setminus\{w\}]$, i.e. $G^{\prime}$ is the graph obtained from $G$ by removing the vertex $w$. $S$ is still a 2-incident independent set in $G^{\prime}$ (but now it may contain twins). Notice that $\delta_{G^{\prime}}(S)=\delta_{G}(S)-1$. The case where $S$ contains no twins contradicts the induction hypothesis. Suppose that $S$ does contain two twins, say $x$ and $y$. Then, their neighbourhood in the original graph $G$ differs only by the vertex $w$ (thus there can only be pairs of twins, e.g. no triplets). Furthermore, $S\setminus\{x,y\}$ is 2-incident and $G^{\prime}\star^{2}S=(G^{\prime}\star^{2}S)\star^{1}\{x\}$. Let $S^{\prime}$ be the set obtained from $S$ by removing every pair of twins and every vertex of degree 1. $S^{\prime}$ is 2-incident and $|S^{\prime}|\leqslant 10$. A non-empty 2-incident independet set without twins or vertex of degree 1 is genuine, and thus by Lemma 3.24 has at least 11 vertices. Thus, $S^{\prime}=\emptyset$.
@@
As $|\delta_{G}(S)|\geqslant 7$, $|N_{G}(x)\cap(\delta(S)\setminus\{u,v\})|\geqslant 5$. For any $w\in\delta(S)\setminus\{u,v\}$, in $G[V\setminus\{w\}]$ each vertex of degree at least 2 has a twin. Thus, in $G$, for any $w\in\delta(S)\setminus\{u,v\}$, there is a vertex $x_{w}\in S$ such that $N_{G}(x_{w})=N_{G}(x)\setminus\{w\}$. Additionally, for any such vertex $x_{w}$, for any $w^{\prime}\in\delta(S)\setminus\{u,v,w\}$, there is a vertex $x_{w,w^{\prime}}\in S$ such that $N_{G}(x_{w,w^{\prime}})=N_{G}(x_{w})\setminus\{w\}$. Thus, $S$ contains at least $161+5+\binom{5}{2}=$ 16 vertices, contradicting the hypothesis that $|S|\leqslant 12$.

### v3

@@ Appendix F Computer-assisted study of 2-local complementation
  • $|V^{\prime}\setminus S^{\prime}|\leqslant|V\setminus\textup{{supp}}(S)|$;
  • there exists no set $A\subseteq S^{\prime}$ such that $G^{\prime}\star^{2}S^{\prime}=G^{\prime}\star^{1}A$.
@@
The proof is constructive, in the sense that we construct $G^{\prime}$ and $S^{\prime}$ from $G$ and $S$. According to Proposition˜2.4, there exists set $S_{1},S_{2}\subseteq\textup{{supp}}(S)$ such that $S_{2}$ is 2-incident and $G\star^{2}S=G\star^{2}S_{2}\star^{1}S_{1}$. Also, the existence of a set $A\subseteq S_{2}$ such that $G\star^{2}S_{2}=G\star^{1}A$ implies $G\star^{2}S=G\star^{1}(A\Delta S_{1})$. Let $G^{\prime}=G$ and $S^{\prime}=S_{2}$, then apply the following operations:
@@
    Remove each vertex of $S^{\prime}$ of degree 0 or 1;
@@
It is easy to generate each graph of $\mathcal{G}_{k}$, although the number of elements in $\mathcal{G}_{k}$ grows double exponentially fast with $k$. $S$ can be defined as a list of words in $\{0,1\}^{k}$ of weight at least 2. More precisely each vertex of $S$ is uniquely associated with a set of $V\setminus S$ of size at least 2, its neighbourhood. Furthermore, the 2-incidence of $S$ implies that $S$ is uniquely determined by the set of its vertices of degree at least 4. Indeed, starting from a set containing only vertices of degree at least 4, the conditions of the form $S\bullet\Lambda_{G}^{K}=0\bmod 2^{r-k-\delta(k)}$" translate into a procedure to find which vertices of degree 3 then 2 need to be added to the set so that $S$ is 2-incident. This proves that there is a bijection between $\mathcal{G}_{k}$ and lists of words in $\{0,1\}^{k}$ of weight at least 4. Thus, the size of $\mathcal{G}_{k}$ is exactly given by the formula
@@
For $k=1,2,3,4,5$ and $6$, the size of $\mathcal{G}_{k}$ is respectively $1,1,1,2,2^{6}=64$, and $2^{22}\sim 4\times 10^{6}$ which is suitable for computation. But, even for $k$ as low as seven, the size of $\mathcal{G}_{7}$ is $2^{64}\sim 2\times 10^{19}$.
@@
According to Lemma˜F.1, this is enough to prove Lemma˜4.1:
@@
Given a set $S$, we define the strict neighbourhood of $S$ in the graph $G$ as the set of vertices not in $S$ that are connected to at least one vertex in $S$: $\delta_{G}(S)=\{v\in V\setminus S\penalty 10000\ |\penalty 10000\ N_{G}(v)\cap S\neq\emptyset\}$. Let us prove, by induction on $|\delta_{G}(S)|$, that for any 2-incident independent set $S$ such that $S$ does not contain any twins and $|S|\leqslant 12$, $G\star^{2}S=G$.
First, by brute force, we check that the property is true whenever $|\delta(S)|\leqslant 6$ (see details in Appendix˜F, code available at [9]).
@@
Let $w\in\delta(S)\setminus\{u,v\}$, and let $G^{\prime}=G[V\setminus\{w\}]$, i.e. $G^{\prime}$ is the graph obtained from $G$ by removing the vertex $w$. $S$ is still a 2-incident independent set in $G^{\prime}$ (but now it may contain twins). Notice that $\delta_{G^{\prime}}(S)=\delta_{G}(S)-1$. The case where $S$ contains no twins contradicts the induction hypothesis. Suppose that $S$ does contain two twins, say $x$ and $y$. Then, their neighbourhood in the original graph $G$ differs only by the vertex $w$ (thus there can only be pairs of twins, e.g. no triplets). Furthermore, $S\setminus\{x,y\}$ is 2-incident and $G^{\prime}\star^{2}S=(G^{\prime}\star^{2}S\setminus\{x,y\})\star^{1}\{x\}$. Let $S^{\prime}$ be the set obtained from $S$ by removing every pair of twins and every vertex of degree 1. $S^{\prime}$ is 2-incident and $|S^{\prime}|\leqslant 10$. A non-empty 2-incident independent set without twins or vertex of degree 1 is genuine, and thus by Lemma˜3.24 has at least 11 vertices. Thus, $S^{\prime}=\emptyset$.
@@
As $|\delta_{G}(S)|\geqslant 7$, $|N_{G}(x)\cap(\delta(S)\setminus\{u,v\})|\geqslant 5$. For any $w\in\delta(S)\setminus\{u,v\}$, in $G[V\setminus\{w\}]$ each vertex of degree at least 2 has a twin. Thus, in $G$, for any $w\in\delta(S)\setminus\{u,v\}$, there is a vertex $x_{w}\in S$ such that $N_{G}(x_{w})=N_{G}(x)\setminus\{w\}$. Additionally, for any such vertex $x_{w}$, for any $w^{\prime}\in\delta(S)\setminus\{u,v,w\}$, there is a vertex $x_{w,w^{\prime}}\in S$ such that $N_{G}(x_{w,w^{\prime}})=N_{G}(x_{w})\setminus\{w\}$. Thus, $S$ contains at least $1+5+\binom{5}{2}=16$ vertices, contradicting the hypothesis that $|S|\leqslant 12$.